1) Use both the ionic and covalent counting systems as was done in the examples in class to give electron counts for the metaPnP .BhP Php -00 H Co Co(m.C.Hg)(CO),

ноос COOH Note: Divalent NR, is an X ligand Pdf ד
Answer #1

Solution:

Br -Br Covalent method: - Valence e- of zo = 4 Electront count Bor Br = 102=2 Electron count = 5x2 for o u = 10- Cas heptacit

Tonic Mothony Election Count Covalent Method tehlico Electron count Co Co=2 O =5 fe = 8 -CH₂ CH₂ = 1 r Total electron Count 2Covalent Method (Electron count) R$ 8-17 -NH₂ - NH2 = 2 Total Electron Count = - +1+2 +6 = 163) Ionic Method: oxidation State

Electron count for (a) to (d) has been founded.

As per Chegg's policy, I've solved first 4 parts.

Hope, it helps. :)